Find the values of x for which f(x) = g(x).
1. f(x) = x2 + 2x + 1, g(x) = 6x − 2
2. f(x) = x4 − 18x2, g(x) = 18x2

Answer by Alan3354(69443)

Find the values of x for which f(x) = g(x).
1. f(x) = x2 + 2x + 1, g(x) = 6x − 2
x^2 + 2x + 1 = 6x - 2
Solve for x.
============
2. f(x) = x4 − 18x2, g(x) = 18x2
Same as #1.